Balance this equation, and then answer the question:

CH3 ​OH + O2 CO2​ + H2​O



 If you combust 56.62 g of CH3OH, how many grams of H2O will be formed?


1
Expert's answer
2021-04-12T03:44:49-0400

2CH3OH+3O2→2 CO2+4H2O

2 moles of CH3OH = 4 moles of H2O

Molar Mass of CH3OH = 32.04

= 56.62/32.04 = 1.7672 moles

= 1.7672 × 2 = 3.5344 moles of H2O

Molar Mass of H2O = 18.01

= 3.5344 × 18.01 = 63.654544 g of H2O